### v2

# Beyond Matryoshka: Revisiting Sparse Coding for Adaptive Representation

Abstract
Many large-scale systems rely on high-quality deep representations (embeddings) to facilitate tasks like retrieval, search, and generative modeling. Matryoshka Representation Learning (MRL) recently emerged as a solution for adaptive embedding lengths, but it requires full model retraining and suffers from noticeable performance degradations at short lengths. In this paper, we show that sparse coding offers a compelling alternative for achieving adaptive representation with minimal overhead and higher fidelity. We propose Contrastive Sparse Representation (CSR), a method that sparsifies pre-trained embeddings into a high-dimensional but selectively activated feature space. By leveraging lightweight autoencoding and task-aware contrastive objectives, CSR preserves semantic quality while allowing flexible, cost-effective inference at different sparsity levels. Extensive experiments on image, text, and multimodal benchmarks demonstrate that CSR consistently outperforms MRL in terms of both accuracy and retrieval speed-often by large margins-while also cutting training time to a fraction of that required by MRL. Our results establish sparse coding as a powerful paradigm for adaptive representation learning in real-world applications where efficiency and fidelity are both paramount. Code is available at this https URL.
Machine Learning, ICML
Figure 1: Overview of our proposed method. (a) Illustrative comparison between standard embeddings (dense, long) and two different compression schemes: Matryoshka representations (MRL) (Kusupati et al., 2022) with short length and our Contrastive Sparse Representation (CSR) based on sparsification. (b) Comparison of retrieval accuracy and time of different methods on ImageNet with GPU. Compared to MRL and int8 quantification (Quant Int8) methods, our sparse embedding approach CSR attains the best retrieval accuracy (very close to full representations) while being much more efficient in retrieval time, using sparse matrix multiplication on GPU. (c) Training GPU hours of CSR compared to baseline methods, where we outperform MRL on 1-NN accuracy with much less training time.
1 Introduction
Representation learning is at the core of deep learning (LeCun et al., 2015) and high-quality representations of inputs (e.g., image, text) empower numerous large-scale systems, including but not limited to search engines, vector databases, and retrieval-augmented generative AI (Lewis et al., 2020). However, the rapid growth in data volume poses significant challenges for latency-sensitive applications. It is thus desirable to develop representations of adaptive inference cost that can best trade-off between accuracy and inference speed.
Recently, a class of methods called Matryoshka Representation Learning (MRL) (Kusupati et al., 2022) has drawn a lot of attention and is now officially supported in the latest OpenAI and Google’s Gemini text embedding APIs  (OpenAI, 2024; Lee et al., 2024b) with millions of users and applications. The idea of MRL is to train an ensemble of representations truncated at different lengths (e.g., from 8 to 2048) through joint multi-task training. However, MRL deviates from standard representation learning and requires full parameter updates to the backbone; the joint training also inevitably sacrifices the quality of representations at a noticeable margin (e.g., 5% drop of top-1 accuracy on ImageNet at full representation length). These limitations render MRL a costly and lossy method for adaptive representation.
In this paper, we revisit sparse coding (Lee et al., 2006) as a much faster, lightweight, and high-fidelity approach to achieve adaptive representation. As illustrated in Figure 1(a), instead of truncating the representation length as in MRL, we leverage sparse vectors and sparse matrix factorization to attain computational efficiency. Specifically, we sparsify a full representation at different levels (characterized by $K$, the number of activated neurons). We find that a few numbers of activated neurons (e.g., $4$ to $16$) can preserve the performance of a much longer dense representation (e.g., $2048$ dimensions). This is in sharp contrast to MRL embeddings whose quality deteriorates a lot at such extremely short lengths ($>$10% drop). Therefore, sparse features using sparse vector formats can be stored efficiently with only a few activated neurons. With the help of sparse matrix factorization (with native GPU support in modern deep learning libraries such as PyTorch) [^1] PyTorch’s native sparse vector library can be found at https://pytorch.org/docs/stable/sparse.html., these sparse embeddings can be used for retrieval tasks at a much higher speed with a complexity order of $\mathcal{O}(K)$, where $K$ is very small. In comparison, MRL requires a longer length of representation (e.g. $256$) to attain similar accuracy (if possible), leading to extra slower inference speed. As shown in Figure 1(b), MRL is inferior to our method in terms of both accuracy and retrieval time by a significant margin.
Another key advantage of sparse features is that they eliminate the need to update the entire network. In contrast, MRL—Kusupati et al. (2022) noted—performs poorly unless full-parameter tuning. However, many existing foundation models, such as the multimodal representations in CLIP (Radford et al., 2021) and the text embeddings in NV-Embed (Lee et al., 2024a), are pre-trained as single representations on massive Internet-scale data. Finetuning these models would be relatively expensive compared to sparse methods. Leveraging recent advances in training sparse autoencoders (SAEs) (Cunningham et al., 2023; Gao et al., 2024), we can train a light-weight 2-layer MLP module for sparsifying pre-trained embeddings within a very short period of time (e.g., half of an hour on ImageNet with a single GPU), which is of orders of magnitude faster than MRL, as shown in Figure 1(c).
These pieces of evidence on accuracy, retrieval time, and training time show that sparse features are strong alternatives to MRL methods for producing high-fidelity and computationally efficient representations with a lightweight module and training cost. Our proposed method, Contrastive Sparse Representation Learning (CSR), combines contrastive retrieval and reconstructive autoencoding objectives to preserve the original feature semantics while better tailing it down to the retrieval tasks. We evaluate CSR on a range of standard embedding benchmarks, from image embedding, text embedding, to multimodal embeddings, and compare it against various state-of-the-art efficient embedding models. Extensive experiments show that CSR consistently outperforms MRL and its variants by significant margins in terms of both accuracy and efficiency. Notably, under the same compute budget, CSR rivals MRL’s performance by 17%, 15%, and 7% on ImageNet classification, MTEB text retrieval, and MS COCO retrieval, respectively. Our main contributions are:
  • We propose sparse coding as an alternative approach to adaptive representation learning and demonstrate its numerous advantages over the MRL approach in terms of fidelity, retrieval cost, and training cost.
  • We introduce an effective learning method for sparse adaptive representation, Contrastive Sparse Representation (CSR) Learning. It combines a task-specific sparse contrastive learning loss with a reconstructive loss to maintain overall embedding quality. This generic design consistently improves performance across different tasks like classification and retrieval.
  • We conduct a detailed analysis of CSR, examining various factors and providing a fair comparison with MRL in terms of retrieval time and accuracy. We further validate CSR’s effectiveness across real-world domains and benchmarks, where it achieves competitive performance against heavily trained state-of-the-art MRL models with significantly lower computational costs. On the inference side, CSR delivers a 69× speedup on ImageNet1k 1-NN tasks without compromising performance compared to quantization-based approaches.
2 Related Work
Adaptive Representation Learning.
Recent research has increasingly focused on learning adaptive representations that cater to multiple downstream tasks with diverse computational requirements. Early efforts explored context-based architectural adaptations (Kim & Cho, 2020), dynamic widths and depths in BERT (Hou et al., 2020), and random layer dropping during training to improve pruning robustness (Fan et al., 2019). More recently, Matryoshka Representation Learning (Kusupati et al., 2022) introduced a novel technique for creating flexible, nested substructures within embeddings, enabling fine-grained control over the trade-off between latency and accuracy. This concept has since been extended to various modalities and applications, including large language models (OpenAI, 2024; Nussbaum et al., 2024; Yu et al., 2024), diffusion models (Gu et al., 2023), recommendation systems (Lai et al., 2024), and multimodal models (Cai et al., 2024; Hu et al., 2024). Other works have further explored token reduction in image and video processing (Yan et al., 2024b; Duggal et al., 2024).
Despite these advances, existing methods often do not fully harness the capabilities of large foundation models, highlighting the need for more effective compression strategies. Our proposed sparse coding methodology addresses this gap by providing a lightweight, plug-and-play solution that can be readily applied on top of any foundation model – significantly reducing computational overhead while preserving representational quality.
Sparse Coding.
Sparse coding serves as a powerful technique for compressing high-dimensional signals and extracting salient features (Wright et al., 2010; Zhang et al., 2015), with learned sparse representations often providing additional computational benefits. Prior work has induced sparsity through modifications to model architectures or training protocols, including modifications to attention mechanisms (Correia et al., 2019), applying $\ell_{1}$-norm penalties to neuron activations (Georgiadis, 2019), and promoting sparse activations in large language models (Mirzadeh et al., 2023; Zhang et al., 2024). However, training state-of-the-art foundation models from scratch under these sparsity constraints has proven challenging (Elhage et al., 2022), limiting their current applicability.
Meanwhile, Sparse Autoencoders have achieved notable success in improving the interpretability of foundation models (Cunningham et al., 2023; Yan et al., 2024a), primarily because they uncover semantic information by mapping high-dimensional data onto lower-dimensional subspaces (Cunningham et al., 2023). Building on these insights – and harnessing the inherent advantages of sparse coding – we investigate how SAEs can be further developed to learn adaptive representations with high efficiency, expanding their applicability to a wider range of tasks.
3 Method
Our proposed framework, Contrastive Sparse Representation learning (CSR), is illustrated in Figure 2. Starting from a pre-trained embedding $v\in\mathbb{R}^{d}$, we project it into a sparse representation space $\mathbb{R}^{h}$, selectively activating the most relevant dimensions for adaptive representation learning. We then regularize this hidden space using a reconstruction-based sparse compression loss (Section 3.2.1). Additionally, with theoretical motivations and guarantees provided by (Wang et al., 2024), we introduce a non-negative contrastive loss to expand model capacity.(Section 3.2.2)
3.1 Preliminaries
Problem Formulation.
For simplicity, we first introduce our framework in the context of a classification task. Let $\mathcal{D}^{N}_{db}=\{{(x_{i},y_{i})}_{i=1}^{N}\}$ be a training dataset of size $N$, where $x_{i}\in\mathcal{X}$ are an input sample and $y_{i}\in\mathcal{Y}^{L}$ are corresponding labels with $L$ classes, We obtain an embedding $v=f(x;\theta_{f}):\mathcal{X}\rightarrow\mathbb{R}^{d}$. We can apply exact $\ell_{2}$-based $k$-nearest neighbor (KNN) search for classification, which has $\mathcal{O}(dN)$ complexity. In practice, KNN often employs high-dimensional embeddings (i.e. $d=4096$) to achieve stronger performance, but at the cost of increased computational latency. Our goal is to learn a more compact representation $v^{\prime}\in\mathbb{R}^{m}$ (where $m\ll d$) that balances accuracy and query latency. This shortened embedding can also benefit other downstream tasks such as retrieval and clustering.
Matryoshka Representation Learning.
MRL simultaneously optimizes embeddings at multiple dimensions, as illustrated in Figure 2, to produce representations of variable size. Specifically, let $\mathcal{M}$ be a set of target embedding sizes. For each $m\in\mathcal{M}$, MRL applies an additional linear classifier to the first $m$ dimensions of the embedding vector, $v_{1:m}\in\mathbb{R}^{m}$. This design ensures each truncated representation is explicitly trained via the final loss. Formally, the MRL objective is:
$\mathcal{L}_{\mathrm{MRL}}=\sum_{m\in\mathcal{M}}c_{m}\mathcal{L}_{\mathrm{CE}}\left({\bm{W}}^{(m)}\cdot f(x_{i};\theta_{f})_{1:m};y_{i}\right)$ (1)
where ${\bm{W}}^{(m)}\in\mathbb{R}^{L\times m}$ is the linear classifier weights corresponding to $v_{1:m}$. Each loss term is scaled by a non-negative coefficient $\left\{c_{m}\geq 0\right\}_{m\in\mathcal{M}}$. The multi-granularity arises from selecting dimensions in $\mathcal{M}$, whose size is constrained to at most $\log(d)$, that is, $\lvert\mathcal{M}\rvert\leq\lfloor\log(d)\rfloor$. For example, Kusupati et al. (2022) choose =$\mathcal{M}=\{8,16,\ldots,1024\}$ as the nesting dimensions.
3.2 Contrastive Sparse Representation Learning
As discussed in Section 1, Equation 1 faces two key constraints: it requires class labels $y_{i}$ and necessitates training the backbone parameters $\theta_{f}$. These limitations reduce its utility in scenarios where web-scale datasets often lack labels and fine-tuning large foundation models is prohibitively expensive. To overcome these constraints while preserving computational efficiency, we employ reconstruction-based sparse coding on top of pre-trained models and incorporate a task-specific loss for generalization ability.
3.2.1 Reconstruction-based sparse compression
Reconstruction-based sparse compression reduces data size by preserving only the most essential components, thereby enabling more efficient storage and computation. Specifically, we adopt sparse autoencoders (SAEs) due to their demonstrated ability to capture semantically rich features from foundation models (Cunningham et al., 2023; Yan et al., 2024a).
Sparse AutoEncoder.
SAEs aim to learn efficient data representations by reconstructing inputs while enforcing sparsity in the latent space. For an input embedding, vector $v:=f(x)\in\mathbb{R}^{d}$, we omit $\theta_{f}$ since SAEs do not optimize the backbone parameters. The encoder and decoder of SAE are defined by:
$z_{k}$ $:=\mathrm{TopK}({\bm{W}}_{\mathrm{enc}}(f(x)-\bm{b}_{\mathrm{pre}})+\bm{b}_{\mathrm{enc}})$ (2)
$\widehat{f(x)}_{k}$ $:={\bm{W}}_{\mathrm{dec}}z_{k}+\bm{b}_{\mathrm{pre}},$
[…]
$\mathcal{L}(k)$ $=\left\|f(x)-\widehat{f(x)}_{k}\right\|_{2}^{2}$ (3)
[…]
$\mathcal{L}_{\mathrm{recon}}=\mathcal{L}(k)+\mathcal{L}(4k)/8+\beta\mathcal{L}_{\mathrm{aux}}$ (4)
where $\mathcal{L}_{\mathrm{aux}}=||e-\hat{e}||^{2}_{2}$, $e=f(x)-\widehat{f(x)}$, and $\hat{e}=W_{\mathrm{dec}}z$ is the reconstruction using the top-$k_{\mathrm{aux}}$ dead latents. By default, we set $k_{\mathrm{aux}}=512$ and $\beta=1/32$, following the setting in (Gao et al., 2024). We also offer dynamic sparsity selection, with $k$ ranging from 8 to 256, to accommodate different tasks across various modalities.
[…]
3.2.2 Contrastive-based sparse compression
Furthermore, we combine task-specific losses to enhance adaptation to diverse downstream tasks. Specifically, we employ contrastive learning strategies for two key reasons: i) Unified loss function-it can integrate supervised tasks by treating intra-class data as positive samples and outer-class data as negative samples, following (Huang et al., 2024), while also being adaptable to unsupervised tasks such as retrieval. ii) Performance improvement-can be achieved by enhancing embedding capabilities through in-batch negative sampling with InfoNCE loss, as demonstrated by  (Carlsson et al., 2021; Zhang et al., 2020; Duan et al., 2024). The loss objective can be formulated as:
$\mathcal{L}_{\mathrm{cl}}=-\frac{1}{\mathcal{B}}\sum_{i=1}^{\mathcal{B}}\log\frac{exp(z_{i}^{T}z_{i})}{exp(z_{i}^{T}z_{i})+\sum_{j\neq i}^{\mathcal{B}}exp(z_{i}^{T}z_{j})}$ (5)
By leveraging the non-negative nature of latent variables $z_{i}$ in sparse autoencoders, Equation 5 can be viewed as a variant of the Non-negative Contrastive Loss (NCL) proposed in (Wang et al., 2024). This interpretation enables us to draw on the theoretical guarantees of NCL, as stated in the following theorem:
Theorem 5.
(Wang et al., 2024) Under certain assumptions, the solution $\phi(x)$ is the unique solution to the NCL objective. As a result, NCL features are identifiable and disentangled.
Theoretically guaranteed by Theorem 5, the model is encouraged to utilize a larger number of latent dimensions to reconstruct the input data. This behavior is empirically demonstrated in Figure 6, where we observe a reduction in “dead” dimensions compared to vanilla SAE approaches.
3.2.3 Training Objective
We synergistically optimize the reconstruction loss $\mathcal{L}_{\mathrm{recon}}$ and non-negative contrastive loss $\mathcal{L}_{\mathrm{ncl}}$. The former preserves semantic information with high fidelity, while the latter refines features for diverse downstream tasks. Together, these losses retain essential information and enhance the model’s discriminative power, leading to improved overall performance. The final loss function is formulated as:
$\mathcal{L}_{\mathrm{total}}=\mathcal{L}_{\mathrm{recon}}+\gamma\mathcal{L}_{\mathrm{ncl}}$ (6)
Here, $\gamma$ is a hyperparameter that balances the two loss components and is set to 1 by default.
4 Empirical Analysis
[…]
4.1 Retrieval Time Comparison with MRL
Experiment Setup.
[…]
Analysis.
[…]
Figure 3: Comparision of retrieval time based on different factors. (a) Fixed-scale scenario (1M database): Both methods achieve performance sweet spots at TopK=16, with CSR exhibiting 2.1× speedup over dense embeddings when sparsity exceeds 80%. (b) Scaling scenario ($h=8192$): CSR exhibits increasingly efficient scalability from 0.5M to 10M, with performance gains accelerating at larger scales. This makes it highly practical for real-world applications involving millions of entries.
4.2 Effect of Input Embedding Dimension $\mathbb{R}^{d}$
Experiment Setup.
We conduct experiments to investigate the relationship between fidelity, input embedding dimension $\mathbb{R}^{d}$, and sparsity under a fixed hidden dimension $\mathbb{R}^{h}$ across different architectures. For ViT backbones, we select ViT-S/16 $(d=384)$ and ViT-L/16 $(d=1024)$, setting the hidden dimension to $h=4096$. For ResNet backbones, we choose Resnet18 $(d=512)$ and Resnet50 $(d=2048)$, with the hidden dimension set to $h=8192$. Other parameters are set to their default values. A more detailed experiment setup is provided in Section E.1.
Analysis.
As shown is Figure 4, our empirical analysis reveals an intriguing pattern: as the input embedding dimension increases, performance degradation diminishes Our model exhibits reduced performance degradation. This observation aligns closely with the Johnson-Lindenstrauss Lemma (Johnson et al., 1984), which posits that mappings from high-dimensional to high-dimensional spaces more effectively preserve structural integrity. This insight is particularly significant, as larger embedding sizes generally encode richer information, thereby achieving better downstream performance. By leveraging these high-dimensional embeddings, our approach more effectively retains essential features and relationships within the data.
Figure 4: CSR achieves higher fidelity as the embedding dimensions of the pre-trained models increase. Fidelity percentages of various architectures (ViT and ResNet backbones) at different sparsity levels reveal a consistent trend.
4.3 Effect of Hidden Representation Dimension $\mathbb{R}^{h}$
Experiment Setup.
To explore the effect of the hidden representation dimension $\mathbb{R}^{h}$ on our model, we use ViT-Large and ResNet50 as pre-trained backbones, following the analysis in Section 4.2. We vary the hidden representation dimension $h$ from $d$ to $16d$ while keeping all other parameters at their default values. Additional implementation details are provided in Section E.2.
Analysis.
Figure 5 compares model performance across different hidden dimensions under varying sparsity constraints. Notably, a shift in the performance trend occurs at $h=4d$. When $h<4d$, performance gradually improves with increasing hidden dimension, reaching its peak at $h=4d$. However, beyond this point, further increases in $h$ lead to performance degradation, particularly under higher sparsity constraints. This trend aligns with the observations of (Gao et al., 2024), which suggest that excessively large hidden dimensions may not be fully utilized, ultimately diminishing model performance. A similar pattern is observed in ResNet. Based on these findings, we set $h=4d$ as the default configuration for all subsequent experiments unless otherwise specified.
Figure 5: CSR exhibits a rise-and-fall trend across different models and hidden dimensions. Model performance across varying hidden dimensions shows that performance peaks at $h=4d$ but degrades beyond this, especially with higher sparsity.
Figure 6: Comparison of dead latent fractions across loss combinations under varying sparsity constraints. Results show that even equipped with $\mathcal{L}_{\text{auxk}}$ and Multiple-TopK at extreme sparsity levels (i.e., k=8,16,32). CSR further alleviates this issue, outperforming baselines and demonstrating its robustness. Evaluations span TopK values from 8 to 1024 using ResNet50 on ImageNet.
4.4 Effect of Different Losses
Experiment Setup.
To investigate the impact of different loss functions on model capacity, particularly in addressing the dead latents problem discussed in Section 3.2.1, we conduct experiments using ResNet50 as the backbone. Following the findings in Section 4.3, we set $h=4d$, while keeping all other parameters at their default values.
Figure 7: (Left & Middle): Results [^3] We use weights from (Kusupati et al., 2022) to measure MRL and MRL-E’s performance.of ImageNet 1-NN accuracy across active dimensions (a) and Normalized Retrieval Time (b). Figure 2 (a) shows that CSR  and SAE outperform all baselines across active dimensions, with CSR  particularly excelling in extreme low-dimensional scenarios (active dim = 8, 16, 32), mitigating performance degradation significantly. Figure 2 (b) highlights CSR ’s superior accuracy-compute trade-off, maintaining pre-trained model performance while achieving high retrieval efficiency through sparse representation. (Right): Task-Specific Evaluation results of CSR-32 and baselines are shown in Figure 2 (c). For each task, the model is trained on three datasets and evaluated on three unseen datasets. The embeddings learned by CSR  effectively preserve the generalization capability of the pre-trained model, demonstrating its robustness and practical applicability across diverse tasks.
Analysis.
[…]
5 Benchmark Results and Analysis
We evaluated the effectiveness of our proposed CSR framework across three mainstream representation modalities: vision, language, and vision+language. For vision representation (see Section 5.1), we conduct image classification on ImageNet-1K and evaluate performance using 1-NN accuracy, following (Kusupati et al., 2022). For language representation (see Section 5.2), we focus on three primary tasks: text classification, text clustering, and text retrieval on the MTEB benchmark (Muennighoff et al., 2022). For multimodal representation (see Section 5.3), we report Recall5 performance on two widely-used datasets: MS COCO (Lin et al., 2014) and Flickr30K (Young et al., 2014). Through these experiments, we aim to provide a holistic understanding of the capabilities of our proposed framework.
5.1 Vision Representation Comparision
Experiment Setup.
We evaluate 1-NN accuracy on ImageNet1k classification, following (Kusupati et al., 2022). The normalized retrieval time is computed through settings in Section 4.1. We select a pre-trained ResNet50 model from (Wightman, 2019) and trained CSR on its encoded embedding of ImageNet1k. For further implementation details, please refer to Section B.
Table 1: Retrieval Efficiency vs. Performance trade-offs under Matched Conditions. We use NV-Embed-V2 as our pre-trained model, and present its performance in the first line of the table in gray. We analyze Dataset-Specific Evaluation results along two key dimensions: (1) Relative Retrieval Time under matched performance and ii) performance under matched retrieval efficiency. Under matched performance, CSR achieves a remarkable 61× speedup, while under matched retrieval efficiency, it improves performance by 15%, demonstrating its superior balance between speed and accuracy. The maximum values are indicated in bold, while the second-highest values are underlined. Relative Retrieval Time is calculated follows the definition in Section 4.
Text Classification Text Clustering Text Retrieval
Active Retrieval Top-1 Acc (%) $\uparrow$ Top-1 Acc (%) $\uparrow$ NDCG@10 (%) $\uparrow$
Method Dim Time MTOPIntent Banking77 TweetSentiment BiorxivP2P BiorxivS2S TwentyNews FiQA2018 NFCorpus SciFACT
NV-Embed-V2 4096 37.6 93.58 92.20 79.73 53.61 49.60 64.82 62.65 43.97 77.93
Stella-en-1.5B-v5 256 2.6 90.45 86.14 76.75 50.81 46.42 60.07 55.59 36.97 77.48
Jina-V3 256 2.8 78.81 84.08 73.81 38.14 34.39 51.96 55.73 36.63 66.63
Nomic-Embed-V1.5 256 2.7 72.47 83.69 59.20 38.19 31.83 48.56 35.00 32.54 68.24
Gecko-Text-Embedding-004-256 256 2.4 77.82 86.01 72.97 36.28 33.09 50.60 55.54 37.81 70.86
OpenAI-Text-Embedding-3-L-256 256 2.8 70.45 83.19 58.98 35.43 33.86 54.24 50.33 37.94 73.10
Arctic-Embed-L-V2.0 256 2.6 67.69 80.99 59.06 34.25 34.07 30.06 44.69 35.02 69.51
M2V-Base-Glove 256 2.4 59.26 72.39 50.02 32.26 22.34 25.38 11.82 23.15 50.66
Jina-V3 64 1.2 68.12 67.98 71.18 36.89 33.57 50.22 44.18 33.66 68.84
Nomic-Embed-V1.5 64 1.6 62.77 80.63 55.23 34.81 44.61 48.06 10.22 18.96 36,55
Potion-Base-2M 64 1.4 42.50 65.17 52.52 25.78 14.94 27.07 32.08 30.72 64.28
SAE (w/ NV-Embed-V2) 32 1.0 87.43 88.11 75.19 51.02 48.68 58.63 49.18 35.14 66.04
CSR (w/ NV-Embed-V2) 32 1.0 89.86 91.02 78.55 53.49 49.13 63.05 57.54 37.06 71.17
Analysis.
Figure 2(a) and (b) illustrate the comparison of learned representation quality through the 1-NN classification accuracy of ResNet50 models trained and evaluated on ImageNet-1K. Reconstruction-based sparse compression methods (CSR & SAE) outperform MRL and MRL-E due to their training flexibility, enabling application top of powerful pre-trained models. These methods also surpass traditional post-hoc compression techniques (e.g., SVD) and linear probes on random features by increasing the overall model total capacity while keeping active dimensions for each sample unchanged, as discussed in Section 1 and Section 3.2.1. This enhanced capability allows CSR to maintain remarkable robustness, with only 2% performance degradation (73.84 vs 75.16). These results highlight that the proposed CSR design can effectively compress pre-trained embeddings while leveraging the natural benefits of sparse matrix multiplication. More detailed experimental results can be found in Section B.4.
5.2 Text Representation Comparision
Experiment Setup.
We assessed CSR on three key tasks from the MTEB benchmark (Muennighoff et al., 2022), testing it across six datasets for each task. In detail, we conduct evaluations in two distinct settings: Dataset-Specific Evaluation, where CSR is trained and tested on different splits of the same dataset to ensure consistency, and Task-Specific Evaluation, where CSR is trained on one dataset and evaluated on unseen datasets within the same task to rigorously assess its generalization capabilities. We choose NV-Embed-V2 (Lee et al., 2024a) as our pre-trained model and present its performance in gray. For further experimental details, please refer to Section C. We refer to CSR-K as a model with the TopK activations and so as SAE.
Analysis
Table 1 demonstrates the performance of CSR and baseline models across multiple tasks and datasets. CSR not only maintains the strong performance of the pre-trained model but also surpasses baselines under varying resource constraints. Notably, it achieves a 61× speedup at matched retrieval efficiency and a 15% performance improvement at matched computational cost, highlighting its superior balance between speed and accuracy. Furthermore, as shown in Figure 2(c), CSR preserves the generalization capabilities of the pre-trained model, ensuring robust performance across diverse downstream tasks. These results underscore the efficacy and versatility of CSR , demonstrating its strong potential for real-world applications.
5.3 MultiModal Representation Comparision
Experiment Setup.
We evaluate our methods on multimodal retrieval tasks using the widely-used MS COCO and Flickr30K datasets, employing the ViT-B-16 backbone. We conduct a MRL fine-tuning on these datasets as baselines, following the standard MRL training paradigm (Kusupati et al., 2022). The performance of our backbone, using the same fine-tuning procedure, is shown in gray. During training, both SAE and CSR leverage a shared sparse embedding layer for images and text. Additional experimental setup and implementation details are provided in Section D.
Analysis.
Table 2 presents the multimodal retrieval task results across different methods. In general, reconstruction-based methods exhibit relatively low performance degradation on both datasets. Compared to the MRL method, CSR achieves average performance gains of 4.6% and 6.8% on image-to-text retrieval, and 9.1% and 6.5% on text-to-image retrieval across the two datasets. SAE experiences more severe performance degradation compared to CSR, which underlines the efficacy of our design in image-text alignment. However, as the sparsity constraint becomes more stringent, the performance gap between CSR and MRL narrows. Upon further investigation, we find that CSR still suffers from the “dead latents” problem even when equipped with advanced mechanisms. Addressing the mitigation of dead latents in the alignment space remains an open challenge, leaving room for future work and study. For a detailed analysis, please refer to Section D.4.
Table 2: MultiModal Retrieval Recall@5 (%) on image-to-text (I2T) and text-to-image (T2I) retrieval across two benchmark datasets, MS COCO and Flickr30K. CSR outperforms ViT-B-16-MRL across all active dimensions and tasks.
MS COCO Flickr30K
Method Active Dim I2T T2I I2T T2I
ViT-B-16 512 74.42 86.47 91.92 97.79
ViT-B-16-MRL 256 67.12 77.53 80.41 89.89
SAE 71.21 82.58 87.76 95.59
CSR 71.41 83.49 87.98 96.79
ViT-B-16-MRL 128 64.19 73.02 77.56 87.80
SAE 64.67 76.70 81.40 91.20
CSR 69.34 81.04 84.05 93.00
ViT-B-16-MRL 64 62.61 72.43 74.22 84.79
SAE 56.30 69.45 70.58 81.30
CSR 62.75 78.10 76.44 88.50
6 Conclusion
In this paper, we introduce Contrastive Sparse Representation Learning (CSR), a generic learning framework offering a high-fidelity and flexible approach to compress embedding, surpassing existing methods like MRL in various tasks and modalities. We believe CSR paves the way for more efficient and flexible representation learning, especially in scenarios constrained by memory, latency or other computational considerations.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Datasets
For Image embedding Experiment:
  • ImageNet-1K (Deng et al., 2009): ImageNet-1K is a large-scale visual database designed to provide researchers with a comprehensive resource for developing and evaluating computer vision models. It contains 1,000 categories, each with a diverse set of images. Specifically, the dataset includes 1,281,167 training images, 50,000 validation images, and 100,000 test images.
For Text embedding Experiment:
Note that, all datasets mentioned below can be found at MTEB (Muennighoff et al., 2022).
  • MTOPIntent (Li et al., 2020): MTOP is a multilingual dataset introduced in 2021. It comprises 100,000 annotated dialogue sentences across six languages and eleven domains. Designed to serve as a benchmark for multilingual task-oriented semantic parsing, this dataset plays a crucial role in advancing technology in this field.
  • Banking77 (Casanueva et al., 2020): Dataset composed of online banking queries annotated with their corresponding intents, consisting of 13,083 customer service queries labeled with 77 intents.
  • TweetSentimentExtraction (Maggie et al., 2020): Dataset from Kag gle competition. Sentiment classification of tweets as neutral, positive or negative.
  • MassiveScenario (FitzGerald et al., 2022): A collection of Amazon Alexa virtual assistant utterances annotated with the associated intent. For each user utterance the label is a theme among 60 scenarios like ’music’, ’weather’, etc. This is a multilingual dataset with 51 available languages.
  • AmazonReviews (McAuley & Leskovec, 2013): A collection of Amazonreviews designed to aid research in multilingual text classification. For each review the label is the score given by their view between 0 and 4 (1-5 stars). This is a multilingual dataset with 6 available languages.
  • Emotion (Saravia et al., 2018): The dataset consists of English Twitter messages categorized into basic emotions, including anger, fear, joy, love, sadness, and surprise.
  • ArxivClusteringS2S, BiorxivClusteringS2S, BiorxivClusteringP2P (Muennighoff et al., 2022): The BioxivS2S dataset is created using public APIs from bioRxiv. For S2S datasets, the input text is simply the title of the paper, while for P2P the input text is the concatenation of the title and the abstract.
  • TwentyNewsgroupsClustering [^4] https://scikit-learn.org/0.19/datasets/twenty_newsgroups.html: Clustering of the 20 Newsgroups dataset, given titles of article the goal is to find the newsgroup (20 in total). Contains 10 splits, each with 20 classes, with each split containing between 1,000 and 10,000 titles.
  • RedditClusteringP2P (Muennighoff et al., 2022): created for MTEB using available data from Reddit posts [^5] https://huggingface.co/datasets/sentence-transformers/reddit-title-body. The task consists of clustering the concatenation of title+post according to their subreddit. It contains 10 splits, with 10 and 100 clusters per split and 1,000 to 100,000 posts.
  • StackExchangeClustering  (Geigle et al., 2021): Clustering of titles from 121 stack exchanges. Clustering of 25 splits, each with 10-50 classes, and each class with 100-1000 sentences.
  • FiQA2018 (Maia et al., 2018): A dataset for aspect-based sentiment analysis and opinion-based question answering in finance.
  • NFCorpus (Boteva et al., 2016): NFCorpus is a full-text English retrieval data set for Medical Information Retrieval. It contains a total of 3,244 natural language queries, with 169,756 automatically extracted relevance judgments for 9,964 medical documents.
  • SciFACT (Wadden et al., 2020): A dataset of 1.4K expert-written claims, paired with evidence-containing abstracts annotated with veracity labels and rationales.
  • Arguana (Wachsmuth et al., 2018b): The dataset consists of debates from idebate.org, collected as of January 30, 2018. Each debate includes the thesis, introductory text, all points and counters, bibliography, and metadata.
  • CQADupStack (Hoogeveen et al., 2015): A benchmark dataset for community question-answering research. It contains threads from twelve StackExchange subforums, annotated with duplicate question information.
  • Quora Question Pairs [^6] https://paperswithcode.com/dataset/quora-question-pairs: A dataset consists of over 400,000 question pairs, and each question pair is annotated with a binary value indicating whether the two questions are paraphrase of each other.
For Multimodal embedding Experiment:
  • MS COCO (Lin et al., 2014): The MS COCO dataset is a large-scale object detection, segmentation, and captioning dataset. It contains images with complex scenes involving multiple objects, each annotated with labels, bounding boxes, and segmentation masks.
  • Flickr30K (Young et al., 2014): The Flickr30k dataset is a collection of images with corresponding textual descriptions. Each image is annotated with multiple captions that describe the scene, objects, and actions depicted.
Appendix B Experiment Detail on Vision Representation.
B.1 Evaluation Metric
We adopt 1-NN as our evaluation metric, implemented using FAISS (Johnson et al., 2019) with exact L2 search, following the setup in (Kusupati et al., 2022). This approach provides an efficient and cost-effective way to evaluate the utility of learned representations for downstream tasks, as 1-NN accuracy requires no additional training. In detail, we use the training set with 1.3M samples as the database and the validation set with 50K samples as the query set.
B.2 Baselines
We select MRL and MRL-E from (Kusupati et al., 2022) as baselines. This work introduces a novel training paradigm that learns representations of varying lengths. MRL-E is an efficient version of MRL, also proposed in (Kusupati et al., 2022).
B.3 Implementation Detail
We select the ResNet50 model [^7] https://huggingface.co/timm/resnet50d.ra4_e3600_r224_in1k as our backbone from (Wightman, 2019). For image preprocessing, we adopt the same procedure as described in (Kusupati et al., 2022; Leclerc et al., 2023). Consistent with (Gao et al., 2024), we utilize a tied encoder-decoder structure to build the CSR framework. The implementation of CSR is based on the codebase [^8] https://github.com/openai/sparse_autoencoder provided by OpenAI. All experiments are conducted on a server equipped with 4 RTX4090 GPUs. The selection of hyperparameters are:
Backbone $d$ $h$ lr epoch Batch Size $k_{\mathrm{aux}}$ $\beta$ $\gamma$ $\mathbb{K}$ Optimizer weight decay eps
Reset50 2048 512 4e-5 10 1024 512 1/32 1.0 8,16,32…2048 Adam 1e-4 6.25 * 1e-10
Table 3: Implementation details on Image experiment.
B.4 1-NN Classification Results
1-NN classification results is shown in B.4. The Int8 quantization operation is performed using a sentence transformer on pre-trained ResNet50 embeddings.
Table 4: 1-NN results of different methods on ImageNet1k classification.
Active Dim MRL MRL-E SVD Rand. LP SAE CSR Quant Int8 Golden Performance
8 62.19 57.45 24.61 6.06 69.69 73.84 - -
16 67.91 67.05 51.02 16.96 71.51 74.39 - -
32 69.46 68.60 64.98 36.99 72.90 74.53 - -
64 70.17 69.61 71.24 51.37 74.12 74.62 - -
128 70.52 70.12 73.63 62.53 74.28 74.78 - -
256 70.62 70.36 74.77 67.32 74.36 74.86 - -
512 70.82 70.74 75.08 72.17 74.38 74.88 - -
1024 70.89 71.07 75.14 74.49 74.40 74.91 - -
2048 70.97 71.21 75.19 75.19 74.41 74.96 73.48 75.19
Appendix C Experiment Detail on Text Representation
C.1 Evaluation Metric
We adopt the universal evaluation metrics used in the MTEB benchmark (Muennighoff et al., 2022). For text classification and clustering, we use top-1 accuracy to assess model performance. For the text retrieval task, we use NDCG@10 (Normalized Discounted Cumulative Gain at 10), a metric that evaluates the quality of a ranked list of items, commonly used in information retrieval and recommendation systems.
C.2 Experiment Setup
We choose three main tasks on MTEB benchmark and randomly select six datasets(for each task) to measure our methods. We also design two experiment settings to evaluate the effectiveness and generalization ability of our methods.
Firstly, we introduce Dataset-Specific Evaluation, where CSR are trained and tested on different splits of the same dataset. We use MTOPIntent (Li et al., 2020), Banking77 (Casanueva et al., 2020) and TweetSentimentExtraction (Maggie et al., 2020) for text classification task. We use BiorxivClusteringS2S, BiorxivClusteringP2P (Muennighoff et al., 2022) and TwentyNewsgroupdClustering for text clustering. For text retrieval, we select FiQA2018 (Maia et al., 2018), NFCorpus (Boteva et al., 2016) and SciFACT (Wadden et al., 2020).
Furthermore, we introduce Task-Specific Evaluation, where CSR are trained and tested on different datasets within the same task to evaluate the generalization ability of our proposed method. We construct a training dataset using the training splits of the aforementioned datasets and test on the corresponding task datasets. For classification: MassivScenario (FitzGerald et al., 2022), AmazonRevies (McAuley & Leskovec, 2013) and Emotion  (Saravia et al., 2018). For clustering: ArxivClusteringS2S, RedditClusteringP2P (Muennighoff et al., 2022) and StackExchangeClustering (Geigle et al., 2021). For retrieval: Arguana (Wachsmuth et al., 2018a), CQADupStack (Hoogeveen et al., 2015) and Quora.
C.3 Baselines
We choose several models that provide MRL embeddings on MTEB benchmark (Muennighoff et al., 2022). These models are Stella-en-1.5B-v5 (Zhang et al., 2025), Jina-V3 (Sturua et al., 2024), Nomic-Embed-V1.5 (Nussbaum et al., 2024), Gecko-Text-Embedding-004-256 (Lee et al., 2024b), OpenAI-Text-Embedding-3-L-256 (OpenAI, 2024), Arctic-Embed-L-V2.0 (Yu et al., 2024), Potion-Base-2M (min, 2024).
C.4 Implementation Detail
We select NV-EmbedV2 (Lee et al., 2024a) as our pre-trained model. We utilize a tied encoder-decoder structure to build the CSR framework. For text classification and clustering tasks, we use data from the same class as positive samples while the other as negative samples to calculate Equation 5. The hyperparameters are set as follows:
Backbone $d$ $h$ lr epoch Batch Size $k_{\mathrm{aux}}$ $\beta$ $\gamma$ $\mathbb{K}$ Optimizer weight decay eps
NV-EmbeV2 4096 16384 4e-5 10 128 1024 0.1 1.0 32,64,256 Adam 1e-4 6.25 * 1e-10
Table 5: Implementation details on Text experiment
Appendix D Experiment Detail on MultiModal Representation
D.1 Evaluation Metric
We adopt the universal evaluation metric Recall@5 to measure performance in the MultiModal Retrieval task. This metric evaluates a model’s ability to retrieve relevant items within its top 5 predictions. Calculated as the fraction of relevant items appearing in the top 5 results out of the total relevant items, a higher Recall@5 indicates better performance in capturing relevant content early in the ranked list, making it useful for recommendation systems and retrieval tasks.
D.2 Experiment Setup
We selected ViT-B-16, trained on the DFN2B dataset [^9] https://huggingface.co/apple/DFN2B-CLIP-ViT-B-16, as our pre-trained model. Following (Kusupati et al., 2022), we implemented MRL in the pre-trained ViT model and fine-tuned it for 50 epochs on the MSCOCO (Lin et al., 2014) and Flickr30K (Young et al., 2014) datasets, respectively. For a fair comparison, we also fine-tuned the backbone on both datasets for 50 epochs using the same hyperparameters, which were then used for the backbone of CSR . The hyperparameters used for fine-tuning are as follows:
Dataset lr epoch Batch Size warmup Optimizer weight decay
MS COCO 5e-6 50 64 10000 Adam 0.1
Flickr30k 5e-6 50 64 10000 Adam 0.1
Table 6: Hyperparameters for fine-tuning ViT-B/16 backbone.
D.3 Implementation Detail
We select the ViT-B-16101010https://huggingface.co/apple/DFN2B-CLIP-ViT-B-16as our backbone from (Wightman, 2019). Consistent with (Gao et al., 2024), we utilize a tied encoder-decoder structure to build the CSR framework. The encoder and decoder structure share between image space and text space. The implementation of CSR is based on the codebase [^11] https://github.com/openai/sparse_autoencoder and OpenCLIP (Cherti et al., 2023). The metric is evaluated through CLIP-benchmark following standard procedure. All experiments are conducted on a server equipped with 4 RTX4090 GPUs. We present detailed training parameters for the multimodal experiment in Table 7.
Dataset $d$ $h$ lr epoch Batch Size $k_{\mathrm{aux}}$ $\beta$ $\gamma$ $\mathbb{K}$ Optimizer weight decay eps
MS COCO 512 2048 4e-4 5 256 512 1/32 1.0 64,128,256 Adam 1e-4 6.25 * 1e-10
Flickr30k 512 2048 4e-4 5 64 1024 1.0 1.0 64,128,256 Adam 1e-4 6.25 * 1e-10
Table 7: Implementation details on MultiModal experiment.
D.4 Disscussion On Dead Latents
Addressing the mitigation of dead latents in the alignment space remains an open challenge, leaving room for future work and study. For a detailed analysis, please refer to Section D.4. Table 2 presents the performance comparison between CSR and MRL, revealing that the gap between the two methods diminishes as sparsity constraints become more stringent. Further analysis indicates that CSR continues to face the “dead latents” issue despite incorporating advanced mechanisms. As shown in Figure 8, CSR exhibits a significant performance drop, corresponding to a sharp rise in dead latent dimensions. We attribute this to a technical challenge, as CSR has demonstrated robust performance in both image and text domains under similar sparsity constraints. This suggests that representations in alignment spaces may require more specialized design, presenting an opportunity for future research and improvement.
Figure 8: Dead latents still exits in image-text alignment space.
Appendix E Empirical Analysis
E.1 Effect on Input Embedding Dimension $\mathbb{R}^{d}$
The implementation details are shown in Table 8. To avoid other unknown factors, we choose ViT-based [^12] https://huggingface.co/timm/vit_small_patch16_224.augreg_in21k_ft_in1k,https://huggingface.co/timm/vit_large_patch16_224.augreg_in21k_ft_in1k and ResNet-based models [^13] https://huggingface.co/timm/resnet18.a1_in1k,https://huggingface.co/timm/resnet50.a1_in1k following same pre-training procedure respectively. To ensure generalizability, we train the model using three different random seeds and report the mean performance in the main paper.
Backbone $d$ $h$ lr epoch Batch Size $k_{\mathrm{aux}}$ $\beta$ $\gamma$ $\mathbb{K}$ Optimizer weight decay eps
ViT-L/16 512 4096 4e-5 10 1024 512 1/32 1.0 8,16,64,256 Adam 1e-4 6.25 * 1e-10
ViT-L/16 1024 4096 4e-5 10 1024 512 1/32 1.0 8,16,64,256 Adam 1e-4 6.25 * 1e-10
ResNet18 512 8192 4e-5 10 1024 512 1/32 1.0 8,16,64,256 Adam 1e-4 6.25 * 1e-10
ResNet50 2048 8192 4e-5 10 1024 512 1/32 1.0 8,16,64,256 Adam 1e-4 6.25 * 1e-10
Table 8: Implementation details on empirical study of input embedding dimension $\mathbb{R}^{d}$
E.2 Effect on Hidden Representation Dimension $\mathbb{R}^{h}$
Implementation details are shown in Table 9. The pre-trained ViT-L/16141414https://huggingface.co/timm/vit_large_patch16_224.augreg_in21k_ft_in1k and ResNet50 models [^15] https://huggingface.co/timm/resnet50.a1_in1k can be found at timm (Wightman, 2019). To ensure generalizability, we train the model using three different random seeds and report the mean performance in the main paper.
Backbone $d$ $h$ lr epoch Batch Size $k_{\mathrm{aux}}$ $\beta$ $\gamma$ $\mathbb{K}$ Optimizer weight decay eps
ViT-L/16 1024 1024 4e-5 10 1024 512 1/32 1.0 8,16,64,256 Adam 1e-4 6.25 * 1e-10
1024 2048 4e-5 10 1024 512 1/32 1.0 8,16,64,256 Adam 1e-4 6.25 * 1e-10
1024 4096 4e-5 10 1024 512 1/32 1.0 8,16,64,256 Adam 1e-4 6.25 * 1e-10
1024 8192 4e-5 10 1024 512 1/32 1.0 8,16,64,256 Adam 1e-4 6.25 * 1e-10
1024 16384 4e-5 10 1024 512 1/32 1.0 8,16,64,256 Adam 1e-4 6.25 * 1e-10
ResNet50 2048 2048 4e-5 10 1024 512 1/32 1.0 8,16,64,256 Adam 1e-4 6.25 * 1e-10
2048 4096 4e-5 10 1024 512 1/32 1.0 8,16,64,256 Adam 1e-4 6.25 * 1e-10
2048 8192 4e-5 10 1024 512 1/32 1.0 8,16,64,256 Adam 1e-4 6.25 * 1e-10
2048 16384 4e-5 10 1024 512 1/32 1.0 8,16,64,256 Adam 1e-4 6.25 * 1e-10
2048 32768 4e-5 10 1024 512 1/32 1.0 8,16,64,256 Adam 1e-4 6.25 * 1e-10
Table 9: Implementation details on empirical study of hidden dimension $\mathbb{R}^{h}$
E.3 Retrieval Time Evaluation
We employ PyTorch (Paszke et al., 2019) to measure retrieval time on ImageNet1k. The average retrieval time is computed over 2000 rounds with a batch size of 512 queries, excluding an initial 100 warm-up rounds. For the learned CSR representation, both query and key embeddings are stored in csr format, and sparse product operations are utilized for similarity computation while maintaining identical experimental settings for fair comparison.